OpenStudy (anonymous):

Determine the x and y intercepts for the line represented by the equation: \ \ 5 x+y+15=0 Part 1: The x-int is of the form (a,0) where a=\ Part 2: The y-int is of the form (0,b) where b=\

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

x-int is when y = 0 so plug in 0 for y then solve for x to get x-int then y-int is when x = 0 so plug in 0 for x then solve for y to get y-int

OpenStudy (anonymous):

I cant get the answer :(

jigglypuff314 (jigglypuff314):

if they gave you x+y+15=0 then x-int is when y = 0 so x + 0 + 15 = 0 solve for x

OpenStudy (anonymous):

is it -15 ?

jigglypuff314 (jigglypuff314):

yep :) then y-int is when x= 0 so 0 + y + 15 = 0 solve for y?

OpenStudy (anonymous):

-15 also right

jigglypuff314 (jigglypuff314):

that's what I got :)

OpenStudy (anonymous):

sorry im bothering you so much cause these questions are due for midnight

jigglypuff314 (jigglypuff314):

I don't mind helping ^_^ but I kinda need to sleep now... sorry :(

OpenStudy (anonymous):

its fine thanks